# CoLo-CAM: Class Activation Mapping for Object Co-Localization in Weakly-Labeled Unconstrained Videos

### v4

@@ Acknowledgment
Appendix A Proposed Approach: Learning with Pixel Pseudo-Labels
In this section, we provide additional details on sampling PLs through the classifier’s CAM ${\bm{C}_{t}}$. These sampled pseudo-labels are used to train the decoder $f$, in combination with other terms. Through the discriminative property of the pseudo-labels, they initiate and stimulate object localization. In CAM-based methods for WSOL, it is common to assume that CAMs’ strong activations are likely to point to the foreground, whereas low activations hint at background regions [22, 87].
In the following, we denote the foreground region as ${\mathbb{C}^{+}_{t}}$ which is computed by the operation ${\mathcal{O}^{+}}$. For simplicity and without adding new hyper-parameters, ${\mathbb{C}^{+}_{t}}$ is defined as the set of pixels with CAM activation ${\bm{C}_{t}}$ greater than a threshold estimated automatically by Otsu method [56] over ${\bm{C}_{t}}$. The background region, i.e., ${\mathbb{C}^{-}_{t}}$, is the remaining set of pixels. Both regions are defined as follows,
@@ Appendix A Proposed Approach: Learning with Pixel Pseudo-Labels
where ${\mathcal{M}(\mathbb{C}^{+}_{t})}$ is a multinomial sampling distribution over the foreground, and ${\mathcal{U}(\mathbb{C}^{-}_{t})}$ is a uniform distribution over the background. This choice of distributions is based on the assumption that the foreground object is often concentrated in one place, whereas the background region is spread across the image. The partially pseudo-labeled mask for the sample ${\bm{X}_{t}}$ is referred to as ${Y_{t}}$, where ${Y_{t}(p)\in\{0,1\}^{2}}$ is the binary label of the location ${p}$ with value 00{0} for the background and 11{1}1 for the foreground. Undefined regions are labeled as unknown. The pseudo-annotation mask ${Y_{t}}$ is then leveraged to train the decoder using partial cross-entropy. Such stochastic pseudo-labels are expected to stimulate the learning of the model and guide it to generalize and respond in the same way over regions with similar color/texture.
Appendix B Additional Visual Results
Fig.6 displays more visual results. Images show that our method yields sharp and less noisy CAMs. However, multi-instances and complex scenes remain a challenge.
Figure 6: Additional localization examples of test sets frames (YTOv1, YTOv2.2). Bounding boxes: ground truth (green), prediction (red). The second column of each method is the predicted CAM on the corresponding image.

### v5

@@ Acknowledgment
Input :  ${\bm{X}_{t}}$: Input frame at time $t$ with its image label $y$, and its corresponding upscaled CAM ${\bm{C}_{t}}$ from a pre-trained classifier.
${f}$: Localization decoder with parameters ${\bm{\theta}}$.
[…]
1 Define foreground region ${\mathbb{C}^{+}_{t}}$ by thresholding the CAM ${\bm{C}_{t}}$ using Ostu method.
2 Define foreground region ${\mathbb{C}^{-}_{t}}$ as the complement of the foreground: ${\mathbb{C}^{-}_{t}=\Omega-\mathbb{C}^{+}_{t}}$ where ${\Omega}$ is the entire image locations.
3 Define ${\mathcal{M}(\mathbb{C}^{+}_{t})}$ as a multinomial sampling distribution over the foreground region. It uses the CAM normalized scores as sampling probabilities.
4 Define ${\mathcal{U}(\mathbb{C}^{-}_{t})}$ as a uniform distribution over the background region.
[…]
6       Sample a pixel location from foreground using ${\mathcal{M}(\mathbb{C}^{+}_{t})}$, and a background pixel using ${\mathcal{U}(\mathbb{C}^{-}_{t})}$.
7       Construct a pseudo-label mask ${Y_{t}}$, where ${Y_{t}(p)\in\{0,1\}^{2}}$, where 00 is for background pixels, and $1$ is for foreground pixels. Pixels that have not been selected in this iteration are set to be unknown.
8       Perform an SGD step over total training loss using Eq.5. The mask ${Y_{t}}$ is used only with partial cross-entropy loss defined in Eq.1.
Algorithm 1 Random pixel samples selection during training illustrated over a single image. Blue background indicates that randomness changes for each frame, and each epoch.
Appendix A Proposed Approach
A.1 Detailed Illustrations for Training and Inference
Fig.6 presents our detailed model for training and inference. During the training, we consider ${n=3}$ input frames. We use three per-frame terms that aim at improving localization at each frame, and without considering any spatiotemporal information. Using the classifier CAM ${\bm{C}_{t}}$, we sample pseudo-labels which are used at the pseudo-label term (PL). Additionally, we apply a CRF, and a size constraint (ASC). In parallel, we apply a multi-frame term ver all the frames allowing explicit communication between them. This is achieved by constraining the output CAMs of the frames to activate similarly over pixels with similar color.
At inference time, we simply consider single frames, i.e., no spatiotemporal information. A bounding box is estimated from the foreground CAM, at the decoder level. In addition, frame classification scores are obtained from the encoder.
Figure 6: Training and inference with our proposed CoLo-CAM method. The single-frame and multi-frame training loss terms are illustrated with $n=3$ frames. Left: training phase: It combines the per-frame terms composed of pseudo-labels loss (PL), absolute size constraint (ASC), and CRF loss, as well as the multi-frame term. Right: inference phase: It operates at a single frame with no temporal dependency, and predicts bounding box localization and classification. Notation and per-frame and multi-frame terms are described in Section 3.
A.2 Proposed Approach: Learning with Pixel Pseudo-Labels
In this section, we provide additional details on sampling PLs through the classifier’s CAM ${\bm{C}_{t}}$. These sampled pseudo-labels are used to train the decoder $f$, in combination with other terms. Through the discriminative property of the pseudo-labels, they initiate and stimulate object localization. In CAM-based methods for WSOL, it is common to assume that CAMs’ strong activations are likely to point to the foreground, whereas low activations hint at background regions [99].
In the following, we denote the foreground region as ${\mathbb{C}^{+}_{t}}$ which is computed by the operation ${\mathcal{O}^{+}}$. For simplicity and without adding new hyper-parameters, ${\mathbb{C}^{+}_{t}}$ is defined as the set of pixels with CAM activation ${\bm{C}_{t}}$ greater than a threshold estimated automatically by Otsu method over ${\bm{C}_{t}}$. The background region, i.e., ${\mathbb{C}^{-}_{t}}$, is the remaining set of pixels. Both regions are defined as follows,
@@ A.2 Proposed Approach: Learning with Pixel Pseudo-Labels
where ${\mathcal{M}(\mathbb{C}^{+}_{t})}$ is a multinomial sampling distribution over the foreground, and ${\mathcal{U}(\mathbb{C}^{-}_{t})}$ is a uniform distribution over the background. This choice of distributions is based on the assumption that the foreground object is often concentrated in one place, whereas the background region is spread across the image. The partially pseudo-labeled mask for the sample ${\bm{X}_{t}}$ is referred to as ${Y_{t}}$, where ${Y_{t}(p)\in\{0,1\}^{2}}$ is the binary label of the location ${p}$ with value 00{0} for the background and 11{1}1 for the foreground. Undefined regions are labeled as unknown. The pseudo-annotation mask ${Y_{t}}$ is then leveraged to train the decoder using partial cross-entropy. Such stochastic pseudo-labels are expected to stimulate the learning of the model and guide it to generalize and respond in the same way over regions with similar color/texture. The sampling process of pixel pseudo-labels is summarized in Algorithm.1.
Appendix B Visual Results.
Fig.7, 8 displays more visual results. Images show that our method yielded sharp and less noisy CAMs. However, multi-instances and complex scenes remain a challenge.
In a separate file, we provide demonstrative videos which show the localization of objects. They highlight additional challenges not visible in still images, in particular:
- Empty frames. Some frames are without objects, yet still show class activations. This issue is the result of the weak global annotation of videos, and the assumption that all frames inherit that same label. It highlights the importance of detecting such frames and treating them accordingly, starting with classifier training. Discarding these frames will help the classifier better discern objects and reduce noisy CAMs. Without any additional supervision, WSVOL is still a difficult task. One possible solution is to leverage the per-class probabilities of the classifier to assess whether an object is present or not in a frame. Such likelihood can be exploited to discard frames or weight loss terms over them.
- Temporal consistency. Although our method brought a quantitative and qualitative improvement, there are some failure cases. In some instances, we observed inconsistencies between consecutive frames. This error is characterized by a large shift in localization where the bounding box (and the CAM activation), moves drastically. This often happens when objects become partially or fully occluded, the appearance of other instances, the scene becomes complex, or zooming in the image makes the object too large. Such setups lead to a sudden shift of the CAM’s focus, leading to an abrupt change of localization which is undesirable. This drawback is mainly inherited from the inference procedure that is achieved at a single frame and without accounting for spatiotemporal information allowing such errors. Future works may consider improving the inference procedure by leveraging spatiotemporal information at the expense of inference time.
Figure 7: Additional localization examples of test sets frames (YTOv1, YTOv2.2). Bounding boxes: ground truth (green), prediction (red). The second column of each method is the predicted CAM over image.
Figure 8: Additional localization examples of test sets frames (YTOv1, YTOv2.2). Bounding boxes: ground truth (green), prediction (red). The second column of each method is the predicted CAM over image.